OpenStudy (anonymous):

find the vertex form of the equation: f(x) = x^2 - x + 5/4 Thanks

OpenStudy (anonymous):

http://prntscr.com/1tzndn click the link it is an image.

OpenStudy (anonymous):

could you draw out the steps, I don't want the answer, I want the process

OpenStudy (tkhunny):

You must "Complete the Square"! f(x) = x^2 - x + 5/4 = (x^2 - x + _____) + 5/4 - ______ The blank is the same both times. What is the magic number that makes the stuff in the parentheses a perfect square trinomial?